OpenStudy (anonymous):

The midpoint of a line segment with end points as (x, 4) and (2, 2) is (2, 1). What is the value of x? x = 6 x = 4 x = 3 x = 2

OpenStudy (anonymous):

Could that @Hero guy help? idk if you can even ask for help here like that..

OpenStudy (anonymous):

never flutter my answer i messed it up lol

OpenStudy (anonymous):

sorry what?

OpenStudy (anonymous):

never mind 0.0

OpenStudy (amistre64):

2 is the center between x and 2 ...

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

so would that make x = 2?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

oh. ok. can you explain why though? i mean is it because it bisects the stuff ro something?

OpenStudy (amistre64):

just as a logical aside: if the distance between the midpoint and endpoint is 0 ... then all points are equal

OpenStudy (anonymous):

im sorry if i come off as like dumb or something, im just horrible at geometry. and ok thanks

OpenStudy (amistre64):

as a formula approach\[\frac{x+2}{2}=2\] \[x+2=4\] etc ...

OpenStudy (anonymous):

wait but wouldnt that make it 4

OpenStudy (anonymous):

oh wait nevermind i got it

OpenStudy (anonymous):

thank you!

OpenStudy (amistre64):

and as i look at it, the midpoint between 4 and 2 cant be 1 on the y side ...

OpenStudy (anonymous):

gotcha

OpenStudy (amistre64):

(x, 4) and (2, 2), mid of (2,1)